29 January 2015 (Thursday)

  • "Legacy Run1 Stripping" waiting on FTS tranfers and recovery from lost files. Large MC campaign upcoming + User jobs but not much happening right now.
  • T0: NTR
  • T1: Any word on an incident report for the lost files at SARA?

26 January 2015 (Monday)

  • "Legacy Run1 Stripping" waiting on FTS tranfers and recovery from lost files. Large MC campaign upcoming + Uer jobs but not much happening right now.
  • T0: NTR
  • T1: NTR

22 January 2015 (Monday)

  • "Legacy Run1 Stripping" campaign "almost" done + MC and user jobs.
  • T0:
  • T1: Ticket opened on Friday to Nikhef/Sara concerning files lost (more than 100K files), closed yesterday: https://ggus.eu/index.php?mode=ticket_info&ticket_id=111205 According to LHCb Datamanagement team 80% of files could not be recovered.

19 January 2015 (Monday)

  • "Legacy Run1 Stripping" campaign "almost" done + MC and user jobs.
  • T0:
  • T1: Ticket opened on Friday to Nikhef/Sara concerning files lost (more than 100K files). Today declared "lost". https://ggus.eu/index.php?mode=ticket_info&ticket_id=111205 (in progress) LHCb Datamanagement team is working on recovering, whatever is possible.

15 January 2015 (Thursday)

  • "Legacy Run1 Stripping" campaign running full steam + MC and user jobs.
  • T0:
  • T1:

12 January 2015 (Monday)

  • "Legacy Run1 Stripping" campaign running full steam and progressing well + MC and user jobs.
  • T0:
  • T1:

8 January 2015 (Thursday)

  • "Legacy Run1 Stripping" campaign running full steam and progressing well, prestaging of input data was restarted + MC and user jobs.
  • T0: Problem of EOS access wednesday between 3h and 5h am. Severe problem with FTS3. FTS tried to transfer file which are not yet staged.(bug in SRM interface of GFAL2, which was discovered and fixed yesterday) and now EOS is returning an unexpected error (in French)
  • T1:
    • RAL : few files lost
    • CNAF : few file lost

5 January 2015 (Monday)

  • "HAPPY NEW YEAR 2015.
  • "Legacy Run1 Stripping" campaign running full steam and progressing well, prestaging of input data was restarted + MC and user jobs.
  • T0:
  • T1:
    • RAL : 1 disk server issue, 1 network glitch during the christmas period

-- JoelClosier - 2015-05-04

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2015-05-04 - JoelClosier
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LHCb All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback